Given a normal distribution with mean equal to 40 and a standard deviation of 6, find the value of x that has 14% of the area to the right.


1
Expert's answer
2021-05-26T16:28:02-0400

"p(x>k=0.14"

i.e "p(x<k)=0.86"

So K correspond to P86

{usng excel}

"=" NORMSINV "(0.86)=1.0803"

Hence Z "=1.0803"

"(K-40)16=1.0803"

"K=40+1.0803\times 6"

"K=46.4819"

"P86=46.4819"

"K=46.48"

So, the value of "x=46.48" that has "14"% of the area to the right